Volume flow rate of water flowing in the pipe Q = 𝑑𝑉/𝑑𝑑 [m^3/ h] varies with time as follows: Q (t) = 5 + 2cos^2 (Ο€t/12). How many cubic meters of water is flowing through the pipe through the day? (the unit of time is an hour)


1
Expert's answer
2020-01-29T05:59:09-0500

Since "\frac{dV}{dt}=Q" , we can calculate V as

"V=\int\limits_0^{24}Q(t)dt=\int\limits_0^{24}(5+2cos^2\frac{\pi t}{12}dt)=\left(5t+t+\frac{6sin\frac{\pi t}{6}}{\pi}\right)_0^{24}=144"